# SAR Strikes Back: A New Hope for RSVQA

### v1

@@ 4.4 Dataset Evaluation
The second limitation, concerns the biases present in the answers. To evaluate the biases in the dataset some scores are introduced in [37], namely: Uniform distribution, Prior distribution, and $L_{B_{score}}$. The uniform distribution is calculated as the inverse of the number of unique answers $A_{\text{unique}}$:
$\text{Uniform}=\frac{1}{A_{\text{unique}}}\quad(1)$
@@
$\text{Prior}=\frac{A_{\text{common}}}{N}\quad(2)$
@@
$L_{B_{\text{score}}}=\frac{\text{Prior}-\text{Uniform}}{1-\text{Uniform}}\quad(3)$
All these scores work in a range of values between 0 and 1. Ideally, a perfect RSVQA dataset would have ${A_{unique}}=\dfrac{N}{A_{common}}$, which means that each answer has the same number of occurrences. In this perfect scenario, the $L_{B_{score}}$ would be equal to 0. In Table 1 the scores are applied to RSVQAxBEN-MM. When the question type is All, it measures the bias by treating the entire dataset as a single group, calculating how much the model favours the most common answer across all questions combined. These scores are also computed by question category, which are then combined on the whole dataset.
We can observe in All and Land Cover in Table 1, that the Uniform distribution has a very low value, which means that there are many different types of answers. Instead in Yes/No the Uniform distribution has a value of 0.50 since there are only two types of answers.
@@
Table 1: Analysis of RSVQAxBEN test set done in [37] using the Prior, Uniform and $L_{B_{score}}$ scores.
@@
5 Performance evaluation
To evaluate our work, we use a VQA metric described in subsection 5.3. In addition, we introduce metrics for classification in subsection 5.1 and subsection 5.2 that we use for the evaluation in the context prediction for Prompt-RSVQA.
5.1 Classification
@@ 6 Experiments
We evaluate our proposed method for RSVQA using optical images (Sentinel-2 RGB) and SAR images (VV and VH polarizations and their ratio). We first train image encoders on the BEN-MM-61 classification task and present the results in subsection 6.1. We compare this encoder with an ImageNet pre-trained encoder for the End-to-end RSVQA pipeline in subsection 6.2. Finally, we present our results obtained with the Prompt-RSVQA pipeline in subsection 6.3.
6.1 Classification
@@
In the first classification experiments, optical and SAR images are used separately. Two networks of different depths, ResNet-50 and ResNet-152 are compared to assess whether greater depth yields better results. The performances of the single-modality encoders are presented in Table 2 (rows a,b,c,d). The performances are evaluated on the classification task using F1-micro and F1-average score. In addition, we assess the performances of the encoders for the RSVQA task using MR and HD.
@@
Based on the single-modality results, we compare the three fusion methods introduced in subsection 3.3 (early fusion, halfway fusion and late fusion) with ResNet-50. The results are presented in the row e,f,g of table  2
@@
In the early fusion, optical and SAR images are concatenated and inserted in a ResNet-50 pre-trained on ImageNet. The first layer of the ResNet is modified to have 6 channels as input instead of 3. In this layer, the weights are initialised using the weights resulting from the optical-only training of the ResNet-50 on BEN-MM-61 for the optical channels and the weights resulting from the SAR-only training of the ResNet-50 on BEN-MM-61 for the SAR channel.
@@
In the halfway fusion, each modality is inserted in a frozen ResNet-50 pre-trained on the BEN-MM-61 classification task, with parameters set as in Table 2[b,d]. The last layer of each network is deleted and the feature maps of the two modalities are concatenated. An average pooling, flattening, linear layer and sigmoid are applied to the concatenated feature maps to finally have the prediction of the classes. This last linear layer is retrained on the classification task with BEN-MM-61.
@@
The late fusion takes as input the two optical and SAR images and inserts them into two different ResNet-50 fully pre-trained on the classification task with the parameters set as in 2[b,d]. The two decision vectors, before being thresholded, are concatenated and inserted in an MLP that produces the final vector, which represents the classes detected in the images. This MLP is also retrained on a classification task with BEN-MM-61.
@@
We train all the models using Adam as an optimizer, a learning rate of $10^{-6}$, a batch size of 64 and a binary cross-entropy as the loss. The number of epochs used to train each model is displayed in Table 2. In this work, all of the models are trained with an Nvidia V100 with 16GB GPU.
6.2 End-to-end RSVQA
@@
In [5], the End-to-End RSVQA was tested only with optical data. In this study, we assess the robustness of the method using SAR images as inputs, and compare it to the results of optical only. For each modality, we compare ResNet-152 and ResNet-50 to extract features compatible with text. First, the networks are frozen and their weights are initialized either with the pre-training on ImageNet or the classification task on BEN-MM-61 as described in section 6.1. In addition, for ResNet-50 initialized on ImageNet only, the network is unfrozen during the full End-to-End pipeline training. The results of the 10 experiments are presented in Table 3
@@
6.3 Prompt-RSVQA
In Prompt-RSVQA, the visual model extracts classes that are transformed in text to help the language model answer the question. Thus, the performances of the Prompt-RSVQA pipeline depend strongly on the classification performances of the visual model. The seven experiments conducted on the classification task are compared in the Prompt-RSVQA pipeline. Their settings are summarized in table  4.
7 Results and discussion
7.1 Classification
By comparing the classification results for optical and SAR images in Table 2[a,b] and Table 2[c,d], we can see that the F1-micro score is 6 to 7% higher using optical images than using SAR images. This highlights the smaller discrimination capacity of SAR images, especially for classes such as green urban areas and wetlands, and in particular inland wetlands.
Table 2[a,b,c,d] also shows that increasing the depth of the neural network, while keeping the same hyper-parameters and the same training time, does not lead to an improvement in classification. The performance of the ResNet-152 network is below the one of ResNet-50 for both modalities. In Table 2[a,b] we can verify that optical results lose 1.6% Indeed, since our classification problem does not require such a complex representation, the additional depth of ResNet-152 may not provide significant benefits. On the contrary, it could worsen the results [38]. Moreover, the greater depth of ResNet-152 could limit its generalization capacities, decreasing its performance since the training set and our test set are geographically separated. This generalization issue could be greater in the case of optical data. Our results are in line with those obtained in [34], in which the shallowest network obtains the best results.
@@ 7.1 Classification
Table 2: Classification results on BEN-MM using ResNet-50/152, using a single modality as input (rows a-d) and both modalities (rows e-g).
* see Section 6.1.
Then, we study the best type of fusion for our task considering only ResNet-50, as it performs the best for each modality individually. From Table 2 we can see that in terms of F1-micro, early fusion has the worst fusion results, followed by halfway fusion. Both have also slightly worse results than the optical image alone. Only late fusion does better than SAR and optical alone, performing the best.
The low results of early fusion are not surprising since it has been shown in [8] that inserting both SAR and optical images in one model may lead to primarily relying on optical channels. The fact that late fusion outperforms halfway fusion is not expected, since we have seen in Section 2 that the relative performances of each fusion depend on the nature of the task, its complexity, as well as on the dataset. Halfway fusion combines features extracted from each domain (optical and SAR) before classification. We believe that for our task, the features extracted from the two domains are not well aligned or containing redundant or conflicting information. The fusion may introduce noise rather than useful information. On the contrary, late fusion combines the predictions of the two modalities rather than their features. This may simplify the integration process since it operates on higher-level abstractions (e.g., class probabilities) rather than complex and potentially incompatible feature spaces. Moreover, combining final predictions might reduce the impact of modality-specific noise. For instance, if one modality is particularly noisy or less informative for certain classes, its influence can be minimized when combined with the other modality’s more accurate predictions. Another advantage of the late fusion mechanism comes from the fact that both classifiers can be designed and trained separately, allowing to take advantage of specific strategies for each modality.
@@
Figure 4 shows the impact of each fusion method on individual L1 classes, reflected by changes in the F1-average scores. These scores account for dataset imbalance and compare the performance of three fusion methods and SAR images with optical images. The model applied to SAR images is generally less accurate for most L1 classes, with the exception of Water bodies. Halfway and late fusion make the most of the improvement of SAR images on the Water bodies, with even larger improvement with respect to SAR modality only. On the other hand, both these fusions are ineffective for the prediction of the Wetlands class, where only the early fusion can improve the prediction accuracy. Also in the Artificial area class, late fusion succeeds in improving only two L3 classes. In contrast, halfway fusion significantly improves the Artificial area class, particularly in three high-frequency classes, whereas early fusion benefits six smaller classes. In the case of Forest and seminatural areas, halfway fusion and late fusion improve almost the same classes, but late fusion has a major impact. Similarly, for Agricultural Areas both early and late fusion improve the performances, but late fusion improves more frequent classes. Our results show that different mergers improve different classes, but what remains common is that if it is not high-frequency classes, the overall performance does not improve significantly.
@@ 7.2 End-to-End RSVQA
Table 3: End-to-End RSVQA results on the RSVQAxBEN-MM dataset.
7.3 Prompt-RSVQA
The results of the Prompt-RSVQA method build on the classification results presented in Section 7.1, even if improvements in classification performance do not directly translate to equivalent gains in VQA accuracy. Indeed, an important observation in classification is that the ResNet-50 model shows a 3.7% difference in MR between optical and SAR predictions. However, this results in a 3.62% increase in Accuracy.
For both SAR and optical images we can observe that the decrease in the F1-micro classification scores when increasing the network depth is reflected in the corresponding VQA results. It is interesting to notice in Table 4[c,d] that with ResNet-152 the land cover accuracy obtained with SAR improves by 0.64%. However, these questions are under-represented in the dataset, which leads to an inferior overall accuracy compared to the one obtained with ResNet-50.
@@ 7.3 Prompt-RSVQA
In the normalized confusion matrix, we observe that for the 30 most-frequent answers the trend is relatively correct, with a visible diagonal. In the non-normalised confusion matrix on the top right corner, the strong imbalance in the answers clearly appears. Indeed, 82% of questions of the test set are yes/no questions. As seen in Table 4[g], our model can predict the answer to these questions with good accuracy. Another fact we can observe is that the model struggles to classify classes that are with the same hierarchy. Two examples of this phenomenon can be illustrated in the blue bounding boxes with answers from the same hierarchy being ordered consecutively in the matrix. Visually, we show in Figure 5 some predictions made by Prompt-RSVQA on the test set.
@@
a OPT ResNet-152 BEN-MM Yes 84.55% 24.93% 73.86% 21.86%
b OPT ResNet-50 BEN-MM Yes 86.07% 26.56% 75.40% 23.40%
c SAR ResNet-152 BEN-MM Yes 82.80% 21.28% 71.77% 19.77%
d SAR ResNet-50 BEN-MM Yes 82.94% 20.64% 71.78% 19.78%
e Early ResNet-50 BEN-MM Yes 85.54% 26.03% 74.88% 22.88%
@@
Table 4: Prompt-RSVQA results on BEN-MM dataset using ResNet with different depth, modality as input and different types of data fusion.
@@
11footnotetext: https://land.copernicus.eu/content/corine-land-cover-nomenclature-guidelines/html/
8 Conclusion
In this article, we introduced a new dataset including SAR images as a modality for RSVQA: RSVQAxBEN-MM. To the best of our knowledge, our work represents the first attempt at using this modality in the RSVQA task. Moreover, we extend two pipelines to use the SAR data as an input. Our experiments with the end-to-end RSVQA model show that it only performs well with optical data, when using ResNet-152 as an image encoder, indicating that it is not a robust model. We show that with Prompt-RSVQA, we can gain in performances with the SAR modality. This suggests that the visual features might be too complex to interact effectively with textual features, especially in networks that are not extensively trained, such as ResNet with ImageNet for other modalities. Furthermore, a model relying on SAR data alone does not achieve results as high as optical data. However, using late fusion, we notice that SAR can add relevant information, particularly in water-related classes. The choice of the fusion method, however, strongly depends on the dataset used. On our proposed dataset, which is imbalanced across classes, late fusion is the most effective method. This result may not hold for other datasets with a different data distribution. Indeed, while early and halfway fusion methods improve the classification results on more classes, these classes are under-represented in questions present in the dataset. As such, future work could be focused on developing new RSVQA datasets including SAR as a modality, with better question coverage (our proposed dataset relates to land cover only) and with improved data distributions.

### v2

@@ IV-D Dataset evaluation)
The second limitation concerns the biases present in the answers. To evaluate the biases in the dataset some scores are introduced in [41], namely: Uniform distribution, Prior distribution, and $L_{B_{score}}$. The uniform distribution is calculated as the inverse of the number of unique answers $A_{\text{unique}}$:
$\text{Uniform}=\frac{1}{A_{\text{unique}}}\quad.$ (1)
@@
$\text{Prior}=\frac{A_{\text{common}}}{N}\quad.$ (2)
@@
$L_{B_{\text{score}}}=\frac{\text{Prior}-\text{Uniform}}{1-\text{Uniform}}\quad.$ (3)
All these scores work in a range of values between 0 and 1. Ideally, a perfect RSVQA dataset would have ${A_{unique}}=\dfrac{N}{A_{common}}$, which means that each answer has the same number of occurrences. In this perfect scenario, the $L_{B_{score}}$ would be equal to 0. In Table I the scores are applied to RSVQAxBEN-MM. When the question type is All, it measures the bias by treating the entire dataset as a single group, calculating how much the model favours the most common answer across all questions combined. These scores are also computed by question category, which are then combined on the whole dataset.
We can observe in All and Land Cover in Table I, that the Uniform distribution has a very low value, which means that there are many different types of answers. Instead in Yes/No the Uniform distribution has a value of 0.50 since there are only two types of answers.
@@
TABLE I: Analysis of RSVQAxBEN test set done in [41] using the Prior, Uniform and $L_{B_{score}}$ scores.
@@
V Performance evaluation
To evaluate our work, we use a VQA metric described in subsection V-C. In addition, we introduce metrics for classification in subsection V-A and subsection V-B that we use for the evaluation in the context prediction for Prompt-RSVQA.
V-A Classification
@@ VI Experiments
We evaluate our proposed method for RSVQA using optical images (Sentinel-2 RGB) and SAR images (VV and VH polarizations and their ratio). We first train image encoders on the BEN-MM-61 classification task and present the results in subsection VI-A. We compare this encoder with an ImageNet pre-trained encoder for the End-to-End RSVQA pipeline in subsection VI-B. Finally, we present our results obtained with the Prompt-RSVQA pipeline in subsection VI-C.
VI-A Classification
@@
In the first classification experiments, optical and SAR images are used separately. Two networks of different depths, ResNet-50 and ResNet-152 are compared to assess whether greater depth yields better results. The last layer of the network is replaced with one that has the correct output number, i.e. 61. The sigmoid is used as an activation function. The performances of the single-modality encoders are presented in Table III (rows a,b,c,d). The performances are evaluated on the classification task using F1-micro and F1-average score. In addition, we assess the performances of the encoders for the RSVQA task using MR and HD.
@@
Based on the single-modality results, we compare the three fusion methods introduced in subsection III-C (early fusion, halfway fusion and late fusion) with ResNet-50. The results are presented in the row e,f,g of Table  III.
@@
In the early fusion, optical and SAR images are concatenated and inserted into a ResNet-50 pre-trained on ImageNet. The first layer of the ResNet is modified to have 6 channels as input instead of 3. In this layer, the weights are initialised using the weights resulting from the optical-only training of the ResNet-50 on BEN-MM-61 for the optical channels and the weights resulting from the SAR-only training of the ResNet-50 on BEN-MM-61 for the SAR channel.
@@
In the halfway fusion, each modality is inserted in a frozen ResNet-50 pre-trained on the BEN-MM-61 classification task, with parameters set as in Table III[b,d]. The last layer of each network is deleted and the feature maps of the two modalities are concatenated. An average pooling, flattening, linear layer and sigmoid are applied to the concatenated feature maps to finally have the prediction of the classes. This last linear layer is retrained on the classification task with BEN-MM-61.
@@
The late fusion takes as input the two optical and SAR images and inserts them into two different frozen ResNet-50 fully pre-trained on the classification task with the parameters set as in III[b,d]. The two decision vectors, before thresholding, are concatenated and fed into an MLP, which generates the final vector representing the detected classes in the images. Considering $n_{c}$ as the number of classes, the two vectors are concatenated after applying the sigmoid function, resulting in an input of size $n_{c}\times 2$ for the MLP. The MLP consists of three layers that progressively transform the vector size. The final output, has a dimention of 61, representing the probabilities of the predicted classes based on the combined information from both modalities. This MLP is also retrained on a classification task using BEN-MM-61.
@@
We train all the models using Adam as an optimizer, a learning rate of $10^{-6}$, a batch size of 64 and a binary cross-entropy as the loss. The number of epochs used to train each model is displayed in Table III. In this work, all of the models are trained with an Nvidia V100 with 16GB GPU. The number of trainable hyperparameters per method is present in Table II.
Model Trainable Parameters
S2/S1 ResNet-152 58’268’797
S2/S1 ResNet-50 23’633’021
Early Fusion 23’642’429
Halfway Fusion 249’917
Late Fusion 67’405
TABLE II: Number of trainable parameters per model.
VI-B End-to-End RSVQA
@@
In [5], the End-to-End RSVQA was tested only with optical data. In this study, we assess the robustness of the method using SAR images as inputs, and compare it to the results of optical only. For each modality, we compare ResNet-152 and ResNet-50 to extract features compatible with text. In both cases, an MLP composed of two layers is added at the end, after the application of a Hyperbolic Tangent function and a dropout of 0.5. First, the networks are frozen and their weights are initialized either with the pre-training on ImageNet or the classification task on BEN-MM-61 as described in section VI-A. In addition, for ResNet-50 initialized on ImageNet only, the network is unfrozen during the full End-to-End pipeline training. The results of the 10 experiments are presented in Table IV.
@@
VI-C Prompt-RSVQA
In Prompt-RSVQA, the visual model extracts classes that are transformed in text to help the language model answer the question. Thus, the performance of the Prompt-RSVQA pipeline depends strongly on that of the visual model. DistilBERT is used as the LLM. After applying the Sigmoid activation function and a dropout of 0.5, two layers are added at the end to match the number of possible answers. The seven experiments conducted on the classification task are compared in the Prompt-RSVQA pipeline. Their settings are summarized in Table  V.
VII Results and discussion
VII-A Classification
By comparing the classification results for optical and SAR images in Table III[a,b] and Table III[c,d], we can see that the F1-micro score is 6 to 7% higher using optical images than using SAR images. This highlights the smaller discrimination capacity of SAR images, especially for classes such as green urban areas and wetlands, and in particular inland wetlands.
Table III[a,b,c,d] also shows that increasing the depth of the neural network, while keeping the same hyper-parameters and the same training time, does not lead to an improvement in classification. The performance of the ResNet-152 network is below the one of ResNet-50 for both modalities. In Table III[a,b] we can verify that optical results lose 1.6% Indeed, since our classification problem does not require such a complex representation, the additional depth of ResNet-152 may not provide significant benefits. On the contrary, it could worsen the results [42]. Moreover, the greater depth of ResNet-152 could limit its generalization capacities, decreasing its performance since the training set and our test set are geographically separated. This generalization issue could be greater in the case of optical data. Our results are in line with those obtained in [37], in which the shallowest network obtains the best results.
@@ VII-A Classification
TABLE III: Classification results on BEN-$M$ using ResNet-50/152, using a single modality as input (rows a-d) and both modalities (rows e-g).
* The network is pretrained on ImageNet, where the first layer is initialized using weights from BEN-MM models (rows b–d). See Section VI-A for details.
Then, we study the best type of fusion for our task considering only ResNet-50, as it performs the best for each modality individually. From Table III we can see that in terms of F1-micro, early fusion has the worst fusion results, followed by halfway fusion. Both have also slightly worse results than the optical image alone. Only late fusion does better than SAR and optical alone, performing the best.
The low results of early fusion are not surprising since it has been shown in [8] that inserting both SAR and optical images in one model may lead to primarily relying on optical channels. The fact that late fusion outperforms halfway fusion is not expected, since we have seen in Section II that the relative performances of each fusion depend on the nature of the task, its complexity, as well as on the dataset. Halfway fusion combines features extracted from each domain (optical and SAR) before classification. We believe that for our task, the features extracted from the two domains are not well aligned or containing redundant or conflicting information. The fusion may introduce noise rather than useful information. On the contrary, late fusion combines the predictions of the two modalities rather than their features. This may simplify the integration process since it operates on higher-level abstractions (e.g., class probabilities) rather than complex and potentially incompatible feature spaces. Interestingly, despite this simpler approach and having the fewest trainable parameters, as shown in Table II, the late fusion strategy still yields the best results among the three fusion schemes. Moreover, combining final predictions might reduce the impact of modality-specific noise. For instance, if one modality is particularly noisy or less informative for certain classes, its influence can be minimized when combined with the other modality’s more accurate predictions. Another advantage of the late fusion mechanism comes from the fact that both classifiers can be designed and trained separately, allowing to take advantage of specific strategies for each modality.
@@
Figure 4 shows the impact of each fusion method on individual L1 classes, reflected by changes in the F1-average scores. These scores account for dataset imbalance and compare the performance of three fusion methods and SAR images with optical images. The model applied to SAR images is generally less accurate for most L1 classes, with the exception of Water bodies. Halfway and late fusion make the most of the improvement of SAR images on the Water bodies, with even larger improvement with respect to SAR modality only. On the other hand, both these fusions are ineffective for the prediction of the Wetlands class, where only the early fusion can improve the prediction accuracy. The low performance of wetlands classification can be attributed to several factors. First, while SAR data, particularly HH polarization, is effective for classifying more structured areas like swamps and uplands, this polarization is not present in the dataset. Additionally, the class ”Swamps” is absent from our dataset, and many of the wetlands that are present, such as fens, bogs, and marshes, exhibit less structured vegetation, leading to misclassification [43]. Moreover, fusing SAR with optical data may introduce inconsistencies due to differences in resolution and spectral properties [44]. To enhance classification accuracy, the use of multispectral Sentinel-2 bands has been shown to improve detection, particularly for smaller wetlands [44]. Similarly, Digital Elevation Models (DEM) have proven useful in distinguishing wetland classes [43]. This aligns with the findings of [8], which demonstrated that adding a ratio channel, and thus incorporating more volumetric information, improved wetland classification by 2%.
In the Artificial area class, late fusion succeeds in improving only two L3 classes. In contrast, halfway fusion significantly improves the Artificial area class, particularly in three high-frequency classes, whereas early fusion benefits six smaller classes. In the case of Forest and seminatural areas, halfway fusion and late fusion improve almost the same classes, but late fusion has a major impact. Similarly, for Agricultural Areas both early and late fusion improve the performances, but late fusion improves more frequent classes. Our results show that different mergers improve different classes, but what remains common is that if it is not high-frequency classes, the overall performance does not improve significantly.
@@ VII-B End-to-End RSVQA
TABLE IV: End-to-End RSVQA results on the RSVQAxBEN-MM dataset.
VII-C Prompt-RSVQA
The results of the Prompt-RSVQA method build on the classification results presented in Section VII-A, even if improvements in classification performance do not directly translate to equivalent gains in VQA accuracy. Indeed, an important observation in classification is that the ResNet-50 model shows a 3.7% difference in MR between optical and SAR predictions. However, this results in a 3.62% increase in Accuracy.
For both SAR and optical images we can observe that the decrease in the F1-micro classification scores when increasing the network depth is reflected in the corresponding VQA results. It is interesting to notice in Table V[c,d] that with ResNet-152 the land cover accuracy obtained with SAR improves by 0.64%. However, these questions are under-represented in the dataset, which leads to an inferior overall accuracy compared to the one obtained with ResNet-50.
@@ VII-C Prompt-RSVQA
Although these relatively simple fusion strategies already provide better performance, it should be noted that more complex approaches have also been explored. For instance, methods such as VisualBERT [45] have been applied in remote sensing VQA settings [46]. However, despite leveraging attention-based fusion mechanisms, the performance reported in that work remains below the level achieved by our simpler architecture.
@@
In the normalized confusion matrix, we observe that for the 30 most-frequent answers the trend is relatively correct, with a visible diagonal. In the non-normalised confusion matrix on the top right corner, the strong imbalance in the answers clearly appears. Indeed, 82% of questions of the test set are yes/no questions. As seen in Table V[g], our model can predict the answer to these questions with good accuracy. Another fact we can observe is that the model struggles to distinguish between classes that belong to the same branch of the CORINE Land Cover hierarchy. Two examples of this phenomenon can be illustrated in the blue bounding boxes with answers from the same hierarchy being ordered consecutively in the matrix. Visually, we show in Figure 5 some predictions made by Prompt-RSVQA on the test set.
@@
a OPT ResNet-152 BEN-$M$ Yes 84.55% 24.93% 73.86% 21.86%
b OPT ResNet-50 BEN-$M$ Yes 86.07% 26.56% 75.40% 23.40%
c SAR ResNet-152 BEN-$M$ Yes 82.80% 21.28% 71.77% 19.77%
d SAR ResNet-50 BEN-$M$ Yes 82.94% 20.64% 71.78% 19.78%
e Early ResNet-50 BEN-$M$ Yes 85.54% 26.03% 74.88% 22.88%
@@
TABLE V: Prompt-RSVQA results on RSVQAxBEN-$M$ dataset using ResNet with different depth, modality as input and different types of data fusion.
@@
VIII Conclusion
In this article, we introduce a new dataset including SAR images as a modality for RSVQA: RSVQAxBEN-MM. To the best of our knowledge, our work represents the first attempt to explore RSVQA for land cover classification using high-resolution SAR imagery, compare its performance with optical images taken under clear conditions and during daylight, and investigate fusion strategies. Moreover, we extend two pipelines to use the SAR data as input. Our experiments with the End-to-End RSVQA model show that it only performs well with optical data when using ResNet-152 as an image encoder, indicating that it is not a robust model. We show that with Prompt-RSVQA, we can gain in performances with the SAR modality. This suggests that visual features may be too complex for effective interaction with textual features in networks that have not been extensively trained for SAR data. Furthermore, a model relying on SAR data alone does not achieve results as high as optical data. However, using late fusion, we notice that SAR can add relevant information, particularly in water-related classes. The choice of the fusion method, however, strongly depends on the dataset used. On our proposed dataset, which is imbalanced across classes, late fusion is the most effective method. This result may not hold for other datasets with a different data distribution. Indeed, while early and halfway fusion methods improve the classification results on more classes, these classes are under-represented in questions present in the dataset. As such, future work could be focused on developing new RSVQA datasets including SAR as a modality, with better question coverage and with improved data distributions. 11footnotetext: https://land.copernicus.eu/content/corine-land-cover-nomenclature-guidelines/html/